Find the equation of the quadratic whose graph is a parabola containing the points (-1,9), (2,15), and (3,29)
Answer by lwsshak3(11628)

Find the equation of the quadratic whose graph is a parabola containing the points (-1,9), (2,15), and (3,29)
****
...............y=ax^2+bx+c
(-1,9)......9=a-b+c
(2,15)....15=4a+2b+c
(3,29)....29=9a+3b+c
..
9=a-b+c
15=4a+2b+c
subtract:
-6=-3a-3b
..
15=4a+2b+c
29=9a+3b+c
subtract:
-14=-5a-b
..
-6=-3a-3b
-14=-5a-b
..
6=3a+3b
42=15a+3b
subtract
-36=-12a
a=3
3b=6-3a=6-9=-3
b=-1
c=9-a+b=9-3-1=5
c=5
..
equation of parabola:
y=3x^2-x+5